\begin{thebibliography}{}


\bibitem{Arn} Arnold, V.I. {\it Geometrical methods in the theory of the ordinary 
differential equations,} second edition. Moscow,  RKhD, 1999.

\bibitem{2} Arnold, V.I.; Ilyashenko, Yu.S. {\it Ordinary differential equations,}
   in: \emph{Dynamical
Systems--1},  Itogi   Nauki  i  Tekhniki,
VINITI, Moscow, 1985; English transl.
in:  \emph{Dynamical Systems~I}, Encyclopaedia Math. Sci.,
  Springer, Berlin, 1988, 1--148.

\bibitem{bjl} Balser, W.; Jurkat, W.B.; Lutz, D.A.  {\it Birkhoff  invariants
and  Stokes'  multipliers  for meromorphic linear  differential
equations,}  J.	Math. Anal. Appl. \textbf{71} (1979), no. 1, 48--94.

\bibitem{buch1} Buchstaber, V.M.; Karpov, O.V.; Tertychnyi, S.I. {\it  	A system on a torus modelling the dynamics of a Josephson junction,} Russian Math. Surveys, 
\textbf{67}:1 (2012), 178--180. 
%
%\bibitem{buch2012}  Buchstaber, V.M.; Karpov, O.V.; Tertychnyi, S.I.  {\it ??????? ?? ????, ???????????? ???????? ???????? ??????????}, ???, \textbf{67}:1 (2012), 165--166.

\bibitem{buch2}  Buchstaber, V.M.; Karpov, O.V.; Tertychnyi, S.I. {\it Rotation number quantization effect }, Theor. and Math. Phys., \textbf{162}:2 (2010), 211--222.  

\bibitem{buch2006}  Buchstaber, V.M.; Karpov, O.V.; Tertychnyi, S.I.  {\it Osobennosti 
dinamiki Dzhozefsonovskogo perekhoda, smeschennogo sinusoidalnym SVCh tokom,} 
Radiotekhnika i elektronika, \textbf{51}:6  (2006), 757--762.

\bibitem{buch2004}  Buchstaber, V.M.; Karpov, O.V.; Tertychnyi, S.I. {\it  	On 
properties of the differential equation describing the dynamics of an overdamped 
Josephson junction}, Russian Math. Surveys,  \textbf{59}:2 (2004), 377--378.

\bibitem{Foote} Foote, R. L., Geometry of the Prytz Planimeter, Reports on Math. Phys, \textbf{42}:1/2, 1998, 249-271.

\bibitem{gi} Gukenheimer, J.; Ilyashenko, Yu. {\it The duck and the devil: canards on 
the staircase,} Moscow Math. J. \textbf{1} (2001), no. 1, 27--47.

\bibitem{LSh2009} Ilyashenko, Yu.S. {\it Lectures of the summer school 
``Dynamical systems'', Poprad (Slovakia), 2009}.

\bibitem{IRF} Ilyashenko, Yu.S.; Ryzhov, D.A.; Filimonov, D.A. {\it Phase-lock effect for equations modeling resistively shunted Josephson junctions and for their perturbations,} 
Functional Analysis and Its Applications  \textbf{45} (2011), no. 3,	192--203.

\bibitem{12} Ilyashenko, Yu.S; Khovanskii, A.G. {\it Galois groups, Stokes operators and a theorem of Ramis,}  Functional Analysis and Its Applications, 
\textbf{24}:4 (1990),  286--296. 

\bibitem{jlp} Jurkat, W.B.; Lutz, D.A.;   Peyerimhoff, A.   {\it Birkhoff
invariants and	effective  calculations	 for meromorphic linear
differential equations,} J. Math. Anal. Appl. \textbf{53} (1976), no. 2, 438--470.

\bibitem{RK} Klimenko, A.; Romaskevich, O. Asymptotic properties of Arnold tongues
and Josephson effect. --- Manuscript.

%\bibitem{Katok} ?. ?. ?????, ?. ??????????. ???????? ? ??????????? ?????? ???????????? ?????? = Introduction to the Modern Theory of Dynamical Systems / ???. ? ????. ?. ????????? ??? ??????? ?. ?????????. ? ?.: ?????????, 1999.

\bibitem{luz} Luzin, N.N. {\it On the approximate integration method of 
academician S.A.Chaplygin,} Uspehi Matematicheskih Nauk, \textbf{6}:6 (46) (1951), 
3--27. 

\bibitem{sib} Sibuya,  Y.  {\it Stokes phenomena,}  Bull.
Amer. Math. Soc., \textbf{83} (1977),  1075--1077.



\end{thebibliography}

\pdfoutput=1 
\documentclass[11pt]{article}
\usepackage{amssymb, latexsym, amsmath, theorem}
\numberwithin{equation}{section}

\theoremstyle{plain}
\theorembodyfont{\itshape}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\theorembodyfont{\rmfamily}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{remark}[theorem]{Remark}
\newenvironment{proof}{{\noindent \textbf{Proof}\,\,}}{\hspace*{\fill}$\Box$\medskip}

\title{On the adjacency quantization in the equation modelling the 
Josephson effect}

\author{Alexey Glutsyuk 
\thanks{Laboratoire
J.-V.Poncelet (UMI 2615 du CNRS and the Independent University of Moscow)} 
\thanks{National Research University Higher School of Economics.}  
\thanks{Permanent address: CNRS, Unit\'e de Math\'ematiques
Pures et Appliqu\'ees, M.R., \'Ecole Normale Sup\'erieure de Lyon,
46 all\'ee d'Italie, 69364 Lyon 07, France. Email: 
aglutsyu@ens-lyon.fr}, \ 
Victor Kleptsyn$^*$\thanks{Permanent address: 
CNRS,  Institut de Recherche Math\'ematique de Rennes 
(UMR 6625 CNRS), Campus de Beaulieu, b\^at. 23, 263 Avenue du G\'en\'eral 
Leclerc, CS 74205, 35042 Rennes Cedex, France. Email: victor.kleptsyn@univ-rennes1.fr}
\ Dmitry Filimonov \thanks{Moscow  Institute of Physics and Technology,  
State University}
\thanks{Moscow State Railway University. Email: mityafil@gmail.com}
\ Ilya Schurov \thanks{National Research University Higher School of Economics. Email: ilya@schurov.com}\ \footnotemark[3]\ \thanks{The present paper uses results  obtained 
 under the support of the project No 11-01-0239 ``Invariant manifolds and asymptotic 
 behavior of slow-fast mappings''  within the program  ``The  HSE scientific foundation''  in 2012/2013}
\thanks{Research of all the authors was  partly supported by  RFBR--CNRS joint 
grant  10-01-93115 NTsNIL\_a and RFBR grant  10-01-00739-a. 
Research of the first author (A.Glutsyuk) was partly supported by 
French ANR grant  ANR-08-JCJC-0130-01}
}
\begin{document}
\maketitle

\centerline{\bf\Large 
This is a brief version with plan of the detailed paper written in Russian}

\bigskip

\begin{abstract} In the present paper we investigate two-parametric family of 
non-autonomous ordinary differential equations on the two-torus that model the 
Josephson effect from superconductivity. We study its rotation number as a function 
of parameters and its {\it Arnold tongues} (also called {\it phase locking domains}): 
the  level sets of the rotation number that have non-empty interior. The Arnold 
tongues of the equation under consideration have many non-typical properties: 
the phase locking happens only for integer values of the rotation 
number~\cite{buch2,IRF}; the boundaries of the tongues are given by 
analytic curves~\cite{buch1,buch2004}, the tongues have zero width at the 
intersection points of the latter curves (this yield the adjacency points). Numerical 
experiments and theoretical  investigations  \cite{buch2006,RK} show 
that each Arnold tongue forms an infinite chain of adjacent domains separated by 
adjacency points and going to infinity in asymptotically vertical direction. 
Recent numerical experiments had also shown that for each Arnold tongue 
all its adjacency points lie on one and the same vertical line with 
the integer abscissa equal to the corresponding rotation number.  In the present paper 
we prove this fact for some open domain of the two-parametric families of 
equations under consideration. In the general case we prove a weaker statement: 
the abscissa of each adjacency point is integer; it has the same sign, as the rotation 
number; its modulus is no greater than that of the rotation number. 
The proof is based on the representation of the differential equations under 
consideration as projectivizations of complex linear differential equations on the 
Riemann sphere, see~\cite{buch2004,Foote,IRF}, and the classical theory 
of complex linear equations.
\end{abstract}

\def\td{\mathbb T^2}
\def\rr{\mathbb R}
\def\zz{\mathbb Z}
\def\cc{\mathbb C}
\def\oc{\overline\cc}
\def\diag{\operatorname{diag}}

\section{Introduction}
\subsection{Main results}
We study two-parametric family of ordinary differential equations on the 
torus   $\td=S^1\times S^1$, 
$S^1=\rr\slash2\pi\zz$, with the coordinates $(x,t)$ of the form  

\begin{equation}\dot x=\frac{dx}{dt}=\nu\sin x + a + s \sin t, \ a,\nu,s\in\rr, \ \nu\neq0.\label{jos}\end{equation}
This family of equations, which we will call  {\it class  J equations} for simplicity, 
models the Josephson effect from superconductivity. The value of the parameter $\nu$ 
is supposed to be fixed and non-zero; the variable parameters are $a$ and $s$. 

The period $2\pi$ flow mapping of the equation is a  diffeomorphism of the space circle 
$S^1=S^1\times\{0\}$ 
$$h_{a,s}\colon S^1\to S^1.$$
In the present paper we study its rotation number  $\rho=\rho(a,s)$ as a function of 
the parameters $a$ and $s$. (Scaling convention: the rotation number equals the 
rotation angle divided by $2\pi$.) 
We will call the $a$-axis horizontal  
(and the $a$- coordinate the {\it abscissa}), and the $s$-axis  vertical (and the 
 $s$- coordinate the {\it ordinate}).
\begin{definition} 
The {\it $r$-th Arnold tongue} is the level set 
$\{(a,s)\mid \rho(a,s)=r\}\subset\rr^2$, provided it has a non-empty interior. 
\end{definition}
 
The rotation number of the system~\eqref{jos} has a physical interpretation: 
this is the average tension for a long time interval. The segments of intersection of 
the  horizontal lines with the Arnold tongues correspond to the Shapiro stairs on the 
Wolt-Amper characteristic. Earlier it was shown that 
\begin{itemize}
	\item the Arnold tongues exist only for integer values of the rotation number~\cite{buch2,LSh2009,IRF};
	\item the boundary of each tongue $\rho=r$ consists of two analytic curves that  
		are graphs of functions denoted $a=g^-_r(s)$ and $a=g^+_r(s)$ 
		(see.~\cite{buch1}; Alexey Klimenko independently observed  that this fact 
		follows  immediately from symmetry\footnote{A class J equation has 
		the symmetry   $(x,t)\mapsto (\pi-x,\pi-t)$. This implies that a 
		 parabolic fixed point of the period flow map  can be only a  
		 fixed point $\pm\frac{\pi}2$ of the symmetry, see \cite{RK}. 
		 A proof of an equivalent statement is contained in  
		 \cite[p. 30]{gi}}  of class J equation, see~\cite{RK}); 
	\item each function $g^-_r(s)$, $g^+_r(s)$ has asymptotics of the $r$-th 
	Bessel  function at infinity (found numerically and  justified on the physical level  in~\cite{buch2006}, proved in~\cite{RK}): 
\begin{equation}
\begin{cases}	
& g^-_r(s)=r-\nu J_{r}(-s/\nu)+o(s^{-1/2})\\
&	g^+_r(s)=r+\nu J_r(-s/\nu)+o(s^{-1/2})
\end{cases};
\label{bessel}
\end{equation}

\item thus, each Arnold tongue is an infinite chain of adjacent bounded domains 
that go to infinity in asymptotically vertical direction; the adjacenty 
points of neighbor domains that do not lie on the horizontal axis  $s=0$ are called simply 
{\it adjacencies.}
\end{itemize}

The following fact was found during numerical experiments.
\medskip

{\bf Experimental fact A.} {\it For every fixed $\nu\neq0$ and every $r\in\mathbb Z$ 
all the adjacencies of the $r$-th Arnold tongue lie on the same vertical line  $a=r$.}
\medskip

The main result of the paper is the following theorem that partly proves the above 
experimental fact. 

\begin{theorem}\label{th} The experimental fact A holds for every fixed 
$\nu\neq0$  with $|\nu|\leq 1$. 
For every fixed  $\nu\neq0$ all the adjacencies have integer abscissas. 
The abscissa of each adjacency has the same sign, as the corresponding rotation 
number, and its module is no greater than that of the rotation number. 
The adjacencies corresponding to zero rotation number are exactly those lying on the 
axis $a=0$.   
\end{theorem}

\begin{corollary} \label{slkr} For every $\nu\neq0$ and $r\in\mathbb Z$ there exists a   
  $M=M(\nu,r)>0$ such that all the adjacencies of the $r$-th Arnold tongue 
with ordinates having modulus greater than  $M$ lie on the line $a=r$.
\end{corollary}

Corollary \ref{slkr} follows from the integrality of the abscissas of the adjacencies   
(Theorem  
\ref{th}) and  the asymptotic formula  (\ref{bessel}) for the boundary of the 
Arnold tongues: the latter formula implies that the points of a given   
 tongue with ordinates large enough  lie in the 1-neighborhood of the line $a=r$.  

\begin{remark} \label{rperem} 
It is known that for every $r\in\mathbb Z\setminus0$ the Arnold tongue $\rho(a,s)=r$ 
intersects the horizontal axis  $s=0$ exactly at the point with the abscissa 
 $\sqrt{r^2+\nu^2}$ (see \cite[Corollary 3]{buch1} and~\cite{IRF}).  This is the 
 point of adjacency of neighbor components of the tongue that we will call 
 {\it queer adjacency.} 
\end{remark}

\subsection{Sketch-proof and plan of the detailed paper}

The proof of Theorem \ref{th} is given in Section 3. It is based on the 
interpretation  of a family of  class J equations as a family of projectivizations 
of linear ordinary differential equations on the Riemann sphere and  classical theory 
of complex linear equatons. The above-mentioned interpretation is given by the 
next proposition stated and proved in Subsection 2.2. 
Its reformulation in some other terms was 
obtained  in  ~\cite{buch2004,Foote,IRF}.

\begin{proposition} \label{prop-change} The variable change 
$p=e^{ix}, \ \tau=e^{it}$ 
reduces the family of class J equations to the following family of Riccati equations:  
\begin{equation}\frac{dp}{d\tau}=\frac1{\tau^2}\left(\nu(1-p^2)\frac{i\tau}2+(a\tau+\frac{is(1-\tau^2)}2)p\right).\label{ricc}\end{equation}
The latter is the projectivization of the following family of linear ordinary differential 
equations via the variable change $p=\frac{z_2}{z_1}$: 
\begin{equation}\dot z=\frac{A(\tau)}{\tau^2}z, \ z=(z_1,z_2)\in\cc^2,  \ 
A(\tau)= \left(\begin{matrix}& 0 & \frac{i\nu\tau}2 \\
& \frac{i\nu\tau}2 & \frac{is}2(1-\tau^2)+a\tau\end{matrix}\right).
\label{lin}\end{equation}
\end{proposition}

 \begin{definition} A nondegenerate linear operator in a vector space is  
 {\it projectively identical (trivial),} if its projectivization is identity as the induced map of the 
 projective space. 
 \end{definition}
 
 The integrality of the abscissas of adjacencies for every $\nu\neq0$ is proved in 
 Subsection 3.1. The linear equations (\ref{lin}) 
 corresponding to class J equations have two 
 irregular nonresonant Poincare rank 1 singular points: 0 and  $\infty$ (see 
 Subsection 2.2).

 \begin{proposition} \label{jos-mon-id} 
 The adjacencies correspond to those parameter values, for which  linear equation (\ref{lin}) has a projectively identical monodromy operator 
 along the positive circuit around zero.
 \end{proposition}
 
 \begin{proof} {\bf (sketch).}
 The adjacencies correspond to those parameter values, for which the period $2\pi$ 
 flow map of class J equation is the identity. This easily follows from  elementary 
 properties of the rotation number (monotonicity in $a$ and continuity in both 
 parameters). The complexification of the latter period flow map is 
 conjugate to the projectivization of the monodromy of linear equation (\ref{lin}), by  
Proposition \ref{prop-change}. This implies Proposition \ref{jos-mon-id}. 
\end{proof}
 
 \begin{lemma} \label{lemon} The monodromy matrix of 
 equation (\ref{lin}) is written in appropriate basis of solutions as a product 
 \begin{equation}
 M=\operatorname{diag}(1,e^{2\pi i a}) S_1S_0,\label{decomp}\end{equation} 
 where the matrices  $S_0$ ($S_1$)
are unipotent and respectively lower (upper) triangular.
 \end{lemma}
 
 \begin{proof} {\bf (sketch).} 
 The classical theory of linear equations at nonresonant irregular singularities 
 \cite{2, 12, bjl,jlp,sib} implies that
  the monodromy matrix of a linear equation at a Poincar\'e rank 1 irregular 
nonresonant singularity  in appropriate basis of solutions  
is equal to the product of three matrices: 

- a diagonal matrix, which is the  
monodromy matrix of the formal normal form of the equation;
 
- two additional unipotent matrices $S_1$ and $S_0$, which are inverse to the so-called 
{\it Stokes matrices.}

The above product decomposition is explicitly given in \cite[p. 35]{12}.

The formal normal form of equation (\ref{lin}) is 
 \begin{equation}\begin{cases} & \dot w_1=0 \\
& \dot w_2=\frac1{\tau^2}(\frac{is}2+a\tau)w_2\end{cases}.\label{norm-jos}\end{equation}
Its monodromy matrix is $\operatorname{diag}(1,e^{2\pi i a})$. This together with the 
above product decomposition proves the lemma.
\end{proof}

\begin{proof} {\bf of the integrality of the abscissas of the adjacencies.} The 
projective triviality of the monodromy of equation (\ref{lin}) is equivalent to the 
triviality of all the operators in the product (\ref{decomp}). This follows from the 
unipotence and  lower (upper) triangularity of the matrices $S_0$ ($S_1$). The 
first operator $\operatorname{diag}(1,e^{2\pi i a})$ is trivial, if and only if 
$a\in\mathbb Z$. Therefore, the abscisses of the adjacencies are integer.
\end{proof}

The experimental fact A  for $|\nu|\leq1$ is proved in Subsection 3.2. An additional 
elementary differential inequality 
(a precision of lemma 4 from \cite{buch1}) 
 shows that for  $|\nu|\leq1$ the complement of the $r$-th Arnold tongue to the  horizontal axis  
 $s=0$ lies strictly between the lines $a=r\pm1$. Thus, all its adjacencies should lie on 
 the  line  $a=r$. 

The general case, when  $\nu$ is arbitrary, is treated in Subsection 3.3. It is easy to 
show that the abscissa of an adjacency has the same sign, as the corresponding 
rotation number.  Using the Argument Principle for appropriate solutions of 
Riccati equations (\ref{ricc}), we  show that the modulus of the abscissa of 
each adjacency is no greater than the modulus of the corresponding 
rotation number. This proves the theorem. 

A preparatory material (rotation number, linear equations, 
Stokes operators and monodromy) is contained in Section 2. 

\section{Acknowledgements}
The authors are grateful to V.M.Buchstaber, E.Ghys, Yu.S.Ilyashenko, A.V.Klimenko 
and O.L.Romaskevich for helpful discussions. 

\end{document}